OpenStudy (anonymous):

find the anti derivative of \[\sin \pi x-2\cos \frac{ x }{ 2 }\]

OpenStudy (anonymous):

integrate it

OpenStudy (anonymous):

well for the first part i got \[-\frac{ \cos \pi x }{ \pi} \] I believe that's correct. but I'm having trouble with the second half

OpenStudy (anonymous):

would I use constant multiple rule and then just integrate cosine?

OpenStudy (anonymous):

+4sin(x/2)

OpenStudy (anonymous):

no minus sorry

OpenStudy (anonymous):

To avoid any confusion , go for x/2 = t... then convert dx into dt...

OpenStudy (anonymous):

Oh I see how it's done now. so you just did \[-2(\frac{ 1 }{ \frac{ 1 }{ 2 } })\sin \frac{ x }{ 2 } \] correct?

OpenStudy (anonymous):

and you got -4sinx/2 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeah, that's like divide the {multiplied of 'x'} and do not change anything for 'added or subtracted'... Like if there's cos(ax+ b)... then divide the (multiplied) a... and do not make changes for 'b' after integrating the cos term...